OpenStudy (anonymous):

please help Part A: Nadia rented a stage platform at $410 for 4 days. If she rents the same stage platform for 8 days, she has to pay a total rent of $770. Write an equation in the standard form to represent the total rent (y) that Nadia has to pay for renting the stage platform for x days. (4 points) Part B: Write the equation obtained in Part A using function notation.(2 points) Part C: Describe the steps to graph the equation obtained above on the coordinate axes. Mention the labels on the axes and the intervals. (4 points)

OpenStudy (anonymous):

@greenlegodude57

OpenStudy (anonymous):

Let me see what I can do..

OpenStudy (anonymous):

ok thanks :)

OpenStudy (anonymous):

Well we have the points (4, 410) and (8, 770).

OpenStudy (anonymous):

So we can find out slope, which is 90.

OpenStudy (anonymous):

I mean -90x + y= 50.

OpenStudy (anonymous):

So -90x - y = 50 would be for A?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

Now we have to write -90x - y = 50 in the form of y = mx + b.

OpenStudy (anonymous):

So that would be y = -90x + 50?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

To graph that you will plot the two points (4, 410) and (8, 770) and draw a line threw them. It will have a slope of -90 and a y-intercept of 50.

OpenStudy (anonymous):

That's for Part C ^^

OpenStudy (anonymous):

to put it as function notation would you change the y to a f(x)? like f(x) = -90x + 50?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

ok thank you! your helping me out a lot today

OpenStudy (anonymous):

Np.